OpenStudy (anonymous):

what is the percent and fraction of 0.24

OpenStudy (anonymous):

put 24 over 100 and simplify

OpenStudy (anonymous):

24/100 = 12/50=

OpenStudy (anonymous):

6/25

OpenStudy (anonymous):

the percent is 24% to find percent, you multiply your decimal but 100.

OpenStudy (anonymous):

can you say that again

OpenStudy (anonymous):

its not 12/50

OpenStudy (anonymous):

how can you find deciaml of 0.24?

OpenStudy (anonymous):

0.24 is a decimal, the fraction is 6/25 and the percent is 24%